OpenStudy (anonymous):

A balloon is being inflated at the rate of 3ft^3/min. How fast is the radius growing when the volume of the balloon is 100ft^3?

OpenStudy (amistre64):

looks like we need to work with a formula for the volume of a sphere.

OpenStudy (anonymous):

Would the usual volume rating work here? \[V= \frac{ 4 }{ 3 }\pi r ^{3}\] Or will we need something different to do this?

OpenStudy (amistre64):

thats a good start, now, the simplest way i know of doing this is to just take an implicit derivative. what do we get?

OpenStudy (anonymous):

Something like? \[4\pi r^{2} \]

OpenStudy (amistre64):

yes, but implicit assumes that V and r are functions of time so the chain rule applies V' = 4pi r^2 r', and we are looking for r' r' = V'/(4pi r^2) we know V' is 3, all we need to do is determine r^2 when V=100

OpenStudy (amistre64):

r = (3V/4pi)^(1/3) r^2 = (3V/4pi)^(2/3) r^2 = (300/4pi)^(2/3)

OpenStudy (amistre64):

soo:\[r'=\frac{3}{4\pi r^2}\] soo:\[\large r'=\frac{3}{4\pi (\frac{300}{4\pi})^{2/3}}\]